North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
J27 Blog

A log of activity concerning the J27.



Tuesday 16th February

Print

Saturday’s working party was Richard De Sadeleer, Neal Woods, Chris H, Mike Bloomfield, Dr Pearson, Angie and me. Richard, Neal and Chris concentrated on the J27 and the rest of us on the K1. Friday and Saturday in particular were very cold – if it got above freezing it did not feel like it!

The work on the locos:

  1. Cylinder bores and big end clearances measured.
  2. Big end brasses examined and white metal tidied up
  3. One valve extracted, rings removed and all components cleaned.
  4. Draw bar pins cleaned and satisfactorily die pen tested
  5. Secondary draw bar cleaned and satisfactorily die pen tested.
  6. Main draw bar cleaned and loaded into company van – to be straightened at Graythorpe Forge.

Unfortunately it was not possible to remove the front buffer beam as had been planned as the loco needs to be moved onto road 4 in the running shed first. It is hoped to remove the buffer beam and the LH cross head, which needs attention this coming weekend.

 

Saturday 30th January 2021

Print

Neal Woods, Doctor Ian Pearson, Paul Taylor and myself in attendance today

Dr Ian took the teams temperatures as usual on signing in.

Paul finished the cleaning the top of the cylinder block in the smoke box, this is now immaculate (I believe he ate his dinner off it) and now ready for not only that part of the annual insurance inspection but also a layer of concrete.

Ian removed the brackets and fittings necessary to remove the Vacuum pipe work and hose from the buffer beam. (Add a three link coupling and we've got something like authentic!..)

He then cracked the nuts on both cylinder covers, and these are now ready to remove.

Neal and I spent the first couple of hours splitting both crossheads. The followers and packings will be removed next week when hopefully the buffer beam, cylinder covers, and pistons will all be removed.

Neal then spent some time addressing the locos left hand Big end webbing and brasses, as these had been problematic when removed. These have been fettled, and measurements of both the Left and Right big ends have been taken by Neal. He also took the opportunity to scrape the white metal as the surface was mottled slightly.

I removed the rear valve covers, this is the steam chest completely exposed now, which will be useful for clearing swarf during the boring process.

All of the motion that needs to removed now has been. This is all sitting for the time being at least next to the loco on 5 Road, clean and labelled.

Last Updated on Saturday, 30 January 2021 23:24
 

22nd and 23rd January 2021

Print

 

65894 working party keeping warm in the running shed at Grosmont on 23 January 2021 - Chris Henwood.

Chris Henwood reports that on Friday 22 January, Robson Hewitson was in attendance, and, during the day, the engine and tender were split. The tender now resides temporarily on No 7 Road in Deviation Shed, and the engine has returned to the back of the running shed on No 5 Road while the motion strip down continues. After Robson had finished the shunt, he managed to remove both crosshead cotters in preparation for them to be split. He also began cleaning the remaining motion to make working within the confines of the frames easier and ultimately more pleasant!

On a very cold Saturday 23rd, Ian Pearson, Neal Woods, Richard De Sadeleer, Paul Taylor and Chris were in attendance and the barrel was lit to provide some warmth (see photo). They found neither our crosshead splitting gear, nor that borrowed from the MPD, was suitable due to the gudgeon pin portion of the gear being oversized. The correct pin is apparently at Hopetown so Neal will bring it down next week. Neal offered to inspect both loco and tender and offered advice on his findings. He then stripped down and lapped in both safety valves. The rear safety valve will be fitted with its new base and valve after the cold exam has been completed.

Paul Taylor cleaned the smokebox and has now completely exposed the top of the cylinder block. Although this will need some final prep, it is essentially ready for paint and concrete. Ian Pearson cleaned all of the washout plugs and most of the mud hole doors. The plugs are now stored in Deviation Shed, and the finished mud hole doors labelled and on the running plate. After consultation with Neal, Richard and Chris started work around the block. Richard began stripping the front end of pipe work fittings, clips and the steam heat valve and connections, as well as the shackle, as part of the preparation for the buffer beam to be removed. Chris split the valve rods from their crossheads and levered them forward, after which Richard manned the large copper hammer and persuaded the valves through the two liners, as Chris manipulated the heads to ensure the rings started in each liner. Both were successfully extracted and are laid with the rest of the motion. A start was made to examine, fettle and assemble the big end brasses ready for measuring next week.

There will be a working party next Saturday, and if you are willing to attend please contact Chris direct. Dependent on availability, there could be another working party mid week: again, please contact Chris if you are able to participate. Remember though that under the Covid controls, numbers are limited to those required for the essential work that has to be done, so please do not turn up on spec as you will be (politely, as ever) asked to leave.

 

14th January 2021

Print

Thursday 14 January saw Robson Hewitson at the MPD, assisted by Andrew Jeffery, working on the J27. The left hand big end connecting rod brasses were removed, and all the components for the con rod assembly labelled up. Work was also started on preparing the right hand side for removal. The planned working party for the Saturday (16th) was postponed in the light of the previous day's weather forecast for heavy snow. In the event of course, none fell from the sky, but better safe than sorry.

Last Updated on Monday, 25 January 2021 09:55
 

9th January 2021

Print

Only Chris Henwood, Robson Hewitson and Richard De Sadeleer were present. Robson prepared the left hand connecting rod for removal, Richard cleaned out the smokebox and then assisted Chris trying to clear some blocked tubes. Chris also removed the safety valves and started work cleaning and inspecting the boiler door holes. Not enough was done however, to prepare the boiler for the planned cold exam on the Monday, so this also was cancelled and will be done when the boiler inspector comes to do the re-arranged Q6 steam test.

Last Updated on Monday, 25 January 2021 10:03
 

10th and 11th November 2020

Print

Chris Henwood was helping last Tuesday (3 November) with the J27 boiler washout at the MPD. Chris Kelley unboxed the loco himself and cleaned up all doors and plugs the day before. Chris K washed the boiler out, while Chris H operated the water pump. After this, Chris H serviced the gauge frame. In the process of cleaning the right hand gauge frame components however, he lost the Stirrup (Bishops hat) in a similar fashion to that which we have all experienced: the immediate disappearance of nuts bolts and tools we are using and may accidentally drop - there one minute, gone the next, every possible location checked - but lost. Barry Neesham kindly offered to make up a new one, featured in the attached photo.

Chris H also attended on Thursday, assisting Chris K with boxing up the boiler. Number 14 plug, a BR 3, needed to be replaced with a size 4. This was tapped out and a new plug fitted. The gauge frame was then to be re assembled. After he had finished the right hand gauge frame, Chris H found the 'lost' Bishops hat (because he wasn't looking for it, of course!). On the Tuesday there had been a mass search in Deviation Shed for our stock of gauge frames in an attempt to locate another Bishops hat, but to no avail. There was also much banter, at Chris H expense, about how we've never needed a spare because no-one's ever lost one before! Now there is a spare of course, which is believed would also fit the Q6, and this has found a home in the gauge glass box on the J27, for the next time Chris H loses one.

65894 gauge frame and new Bishop's hat, November 2020 - Chris Henwood

A steam test on the Friday was partly successful in that the locomotive passed its functional brake tests. However, it failed on a top door passing steam despite being nipped. This has since been addressed by the boilershop staff, though it has been commented that remedial work will have to be done to the offending door in the near future. The loco is now stopped of course, during this period of Lockdown 2.

Because of the pandemic and its consequences for operations on the NYMR, J27 working parties have been few and far between this year. As winter approaches however, regular parties would normally start again, but it is hard to make work plans at moment as its future requirements from the NYMR inevitably remain uncertain. But, importantly, what we have is a working locomotive earning its keep (an unexpected but very welcome £7,360 so far this year), and is part of the Railways' pool of locos available for whatever services it is able to run over Christmas and into the New Year. As a result, Barney has suggested that the bottom end of the loco is not touched for the moment, so it can be used in December. Once the two Standard 4s start entering traffic, it will then be possible to start looking at the motion. To maintain availablity therefore, the required winter work will have to be staggered while hopefully still making significant progress. To make a start, Neal Woods and Chris H went to the running shed on Saturday (7 November), and completed most of the backhead valve exam. They are intending to go in again on 21 November and if anyone is willing to join them for the day, please contact Chris Henwood direct.

Last Updated on Monday, 25 January 2021 09:54
 

Friday 31st July 2020

Print

Chris Henwood reports that the outstanding work on the J27 has been successfully completed. All being well the J27 will be off Grosmont light engine tomorrow, and hauling the Pickering/Levisham shuttles all day, being both Yorkshire Day and the first day of NYMR operation.

He had a working party of sorts yesterday, with Chris, Ian Pearson and Andrew Jeffrey signing in for NELPG, and assistance from Robson Hewitson from the NYMR. Close of play saw the connecting rod bush fitted and fettled, and the rod itself fitted, as well as the removal, resealing and re fitting of a few valves that were leaking in the steam test the previous week. As a result, a warming fire was lit for a steam test today (Friday). That was a success. The little end was stone cold to the touch after a fair bit of running up and down the headshunt, and the valves were leak free. The locomotive was disposed just after lunch, and shunted onto Five Road in readiness for tomorrow's service.

Chris is conscious that in spite of some work being required on the locomotive, he has not advertised for anyone to be rostered for J27 working parties. He hopes that working members will understand that following the steam test for the 'shadow running' week, it was found that the right hand connecting bush was defective. As a result, he arranged with Ian Pearson to go in and drop the rod out ready for a new bush to be machined. This was effectively a two man job both for work reasons but, ultimately, Covid 19 security and social distancing. This job was subject to various shunts through the day as well, so it was felt best not to invite a working party along, especially if they were travelling far, to be waiting around, and potentially not get to do much anyway. The NYMR had also made it clear that it wanted the minimum number of people attending working parties, sufficient for the work in hand and no more. Yesterday's working party (Thursday) was very much last minute, as the new bush was only completed on Tuesday. Again, the work involved in fitting the rod required close-ish working, and, along with the other jobs as mentioned, was all done fairly quickly.

Consequently, Chris rightly felt it was best to have the lads who live literally in Grosmont village, and free to attend at short notice, for this particular work.

However, for those missing their J27 fix, the locomotive should be on washout from approximately 17 August , so he is hoping to get an NELPG working party together to unbox the locomotive and do the wash out, with the MPD staff numbers being so reduced because of furloughing.

 

Friday 17th July 2020

Print

65894 stands under the coaling plant at Grosmont on 16 July 2020 - Chris Henwood

As some will be aware, the J27 was due to have a steam test last Friday,17 July, following the fitting of a new right hand injector steam valve. Chris Henwood reports that the J27 was shunted and had a warming fire lit on the Thursday as planned. The locomotive performed satisfactorily in its steam test, but a suspicious 'knock' from the bottom end was picked up as the locomotive was on its way to the pit to be disposed.This turned out to be due to excessive side to side movement on the little end of the right hand connecting rod, and closer inspection revealed a loose bush in the rod. Its planned running in turn for today was therefore postponed.

Yesterday, the connecting rod was disconnected and lowered out of the frames, revealing the bush was well and truly free of the rod. Measurements were taken of the crosshead, rod end and little end pin as a reference for the repair. The MPD has sourced suitable material for a new bush and this will hopefully be turned and push fitted this week. Following this, Chris will arrange a small working party to refit the rod: he will contact direct those required. The running in turn will then be rescheduled, following which the locomotive should be available for traffic.

In the working party report of 3 July, there was reference to the rods being cleaned. Subsequently, following consultation with Richard Pearson, it was decided to paint the rods red. There is photographic evidence that 65894's rods were painted red during its BR days in the early 1960s at York, and it was felt that the paint would ultimately provide protection from further corrosion and pitting, which it has suffered from for two seasons running. A photograph is attached of the J27 under the Grosmont coaling plant on 16 July 2020 as it now appears.

Last Updated on Tuesday, 21 July 2020 19:52
 

Thursday 2nd July 2020

Print

On Thursday, Ian was back again, along with Andrew Jeffrey, Robson Hewitson and Chris Henwood, focussing on the J27. On arrival Andy Wiffin was already busy fitting the water pipe and drain into the workshop, and got finished about 3 pm.

On the J72, Chris had repaired one of the sliding window frames, and managed to refit the glass. Andrew and Ian removed the temporarily fitted window runners, and set about cleaning the interior of the cab around the windows, as well as cleaning up and preparing the fixings. The windows themselves were then fitted (see photo of the now 'finished' cab minus the right hand injector steam valve)

65894 cab interior minus right hand injector steam valve in Deviation Shed on 2 July 2020 - Chris Henwood.

Robson set about cleaning from the off, and achieved an incredibly thorough clean of the boiler barrel and tender, as well as a good go over the inside motion. Detailing of the builders plates and smoke box brass collar have also been completed - the locomotive looks fantastic. Andrew had brought along a set of Overhead Lines warning stickers, which have not been present since the locomotive entered traffic. These were applied and varnished. He then moved on to cleaning the locomotives coupling rods (see photo). Reassembly of the spark arrestor was also started, for which a few components have been fabricated to help tidy up the original setup.

65894 being cleaned in Deviation Shed by Robson Hewitson and Andrew Jeffrey on 2 July 2020 - Chris Henwood.

Please note that No.8 roller shutter door now has a new padlock issued by the MPD. When access to that Road is required, the key will have to be obtained from the MPD when required. We have also had confirmation from Martyn Cannings that, following his inspection last year, the broken rail on No 8 Road is not considered an urgent problem for the limited amount of shunting that takes place there. Its replacement is not therefore being pursued for the moment.

Last Updated on Thursday, 09 July 2020 13:10
 

Saturday 29th February 2020

Print

In the absence of Ian Pearson, Chris Henwood has reported a very busy day on shed at Grosmont yesterday, which made for an interesting working party and a brilliant atmosphere. The presence of sizeable working parties on 44806 (on the pit) and Lucie, consisting of people of varying footplate grades as well as shed staff, meant that there was no shortage of shunter drivers, passed shunters, and senior shed staff for advice and opinion.

In Deviation Shed, there was just Neal Woods and Chris working on the J27. No work was done on the T2/Q6. The J27 was washed out earlier in the week, and its new wheel position, having returned to Deviation, meant the pistons were in a good position to allow the cylinders' inspection and for measurements to be taken. The bores appear to have suffered a slight lack of lubrication. After some physical investigation and testing, and discussions  throughout the day with Barney and colleagues, it appears previous adjustment in the mechanical lubricator deliveries may have been the cause, but also the answer. Work will continue in the shed to assess, adjust and test the system in the coming weeks. Neal completed the necessary work on the R/H valve, including cleaning the face of the cover, meaning both cylinder covers and the valve covers can be fitted on the arrival of the copper gaskets.

Chris spent the morning thoroughly cleaning the firebox and oiling the tube plate. This job is now finished. After lunch the J27 was shunted to No 1 Road pit, and the contents of the ash pan disposed. The grate and ash pan were swilled out with a hose. The J27 was then returned to No 7 Road.  The fusible plugs were also removed, and a quick 'end of day job' saw the fitting of the whistle.

Next steps are to arrange for the cold examination of the boiler by the boiler inspector. Once the NYMR has finished the work on the safety valves, the J27 will not be far away from being boxed up ready for a hot exam after the boiler inspector has been.

 


Page 3 of 9